Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labour-hours, and its standard costs per unit are as follows: Direct materials: 5 kg at $9.00 per kg $ 45.00 Direct labour: 3 hours at $14 per hour 42.00 Variable overhead: 3 hours at $9 per hour 27.00 Total standard cost per unit $ 114.00 The company planned to produce and sell 20,000 units in March. However, during March the company actually produced and sold 24,800 units and incurred the following costs: Purchased 155,000 kg of raw materials at a cost of $7.20 per kg. All of this material was used in production. Direct labour: 65,000 hours at a rate of $15 per hour. Total variable manufacturing overhead for the month was $612,300. 2. What is the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)
Answer #1

Formula for Calculation of Materials Quantity Variance = (Standard Quantity of Material for Actual Output - Actual Quantity Used) * Standard Material Rate

Standard Quantity of Material for Actual Output = Actual Output * Standard Quantity of Material for One Unit of Production

Actual Output = 24,800 Units

Standard Quantity of Material for One Unit = 5 Kg

Standard Quantity of Material for Actual Output = 24,800 * 5

Standard Quantity of Material for Actual Output = 124,000 Kgs

Standard Material Rate per KG = $ 9

Actual Quantity of Material Used = 155,000 Kgs

Material Quantity Variance for March = (124,000 - 155,000) * 9

Material Quantity Variance = (31,000) * 9

Material Quantity Variance = ($ 279,000)

Material Quantity Variance = $ 279,000 Unfavorable Variance